# The possible adelic indices for elliptic curves admitting a rational cyclic isogeny

### v1

@@ 1. Introduction
Theorem 1 (Serre’s Open Image Theorem, [Ser72]).
@@
Serre’s Open Image Theorem is equivalent to the existence of a positive integer $n$ such that $G_{E}=\pi_{n}^{-1}(G_{E}(n))$; the least such $n$ is called the adelic level of $E$. Since Serre proved Theorem 1, there has been substantial progress toward understanding Galois images of non-CM elliptic curves over $\mathbb{Q}$. Notably, Zywina [zywina2015possibleimagesmodell] and Sutherland [MR3482279] conjecturally classified all possible mod $\ell$ Galois images and gave efficient algorithms for computing them. Rouse–Zureick-Brown [MR3500996] and Rouse–Sutherland–Zureick-Brown [MR4468989] did similarly for $\ell$-adic images. In 2022, Zywina [ZywinaOpen] gave a breakthrough algorithm for computing adelic images of elliptic curves.
An important invariant related to the adelic Galois image of a non-CM elliptic curve $E$ is the adelic index, the quantity $\operatorname{GL}_{2}(\widehat{\mathbb{Z}}):G_{E}]$. In general, the adelic index cannot be determined from the $\ell$-adic indices alone because of possible entanglements between the $\ell$-adic images. For example, consider the elliptic curve 37.a1 in the LMFDB. For all primes $\ell$, its $\ell$-adic image is $\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$, yet its adelic index is 2. In fact, the adelic index of any non-CM elliptic curve over $\mathbb{Q}$ must be at least $2$, as noted in [Ser72, Proposition 22]. Moreover, Jones [MR2563740] proved that, when ordered by naive height, almost all elliptic curves over $\mathbb{Q}$ have adelic index $2$. Adelic indices greater than 2 occur for infinitely many $\overline{\mathbb{Q}}$-isomorphism classes [MR2563740, Theorem 4], but these are sparse asymptotically.
Zywina conjectured that there are only finitely many possible adelic indices for non-CM elliptic curves over $\mathbb{Q}$, and he even proposed a specific finite set $\mathcal{I}$ of all adelic indices [zywina2022possibleindicesgaloisimage, ZywinaOpen].
Conjecture 2 (Conjecture 1.5, [ZywinaOpen]).
@@
This conjecture has been verified for all non-CM elliptic curves over $\mathbb{Q}$ currently appearing in the $L$-Functions and Modular Forms Database (LMFDB) [lmfdb], consisting of more than three million curves. The conjecture is closely connected to Serre’s uniformity question, which asks whether there exists a prime $p$ such that for any prime $\ell>p$, the mod $\ell$ Galois representation of any non-CM elliptic curve $E/\mathbb{Q}$ is surjective. It is now widely conjectured that Serre’s uniformity question has an affirmative answer with $p=37$.
Over the past several decades, significant progress toward Serre’s uniformity question has come from the study of rational points on modular curves along Dickson’s classification: if the mod $\ell$ representation is not surjective, its image is contained in a Borel subgroup, the normalizer of a split or nonsplit Cartan subgroup, or an exceptional subgroup. Serre [Ser72] proved that the exceptional case cannot occur for $\ell>13$. Mazur [MR482230] proved that if $\ell>37$, then $G_{E}(\ell)$ is not contained in a Borel subgroup. Bilu and Parent [MR2753610] proved that $G_{E}(\ell)$ cannot be contained in the normalizer of a split Cartan subgroup for sufficiently large $\ell$. Bilu, Parent, and Rebolledo [MR3137477] later showed that the same result holds for every prime $\ell\geq 11$, except possibly $\ell=13$. The latter case was subsequently resolved via quadratic Chabauty by Balakrishnan, Dogra, Müller, Tuitman, and Vonk [MR3961086]. As for the case of nonsplit Cartan subgroups, recent work of Furio and Lombardo [FurioLombardo23] shows that if $\ell>37$ is nonsurjective, then $G_{E}(\ell)$ is precisely equal to the normalizer of a nonsplit Cartan subgroup. Showing that this does not happen for sufficiently large $\ell$ is the remaining problem for the uniformity question, but eliminating this case appears daunting.
Nevertheless, Serre’s uniformity question is known to have a positive answer for certain families of non-CM elliptic curves over $\mathbb{Q}$ due to work of Lemos. He proved that uniformity holds with $p=37$ for the family of non-CM elliptic curves $E/\mathbb{Q}$ for which there exists a prime $q$ such that $G_{E}(q)$ is contained in the normalizer of a split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ [MR3968924]. Further, he proved that uniformity holds with $p=37$ for the family of all non-CM elliptic curves admitting a nontrivial rational cyclic isogeny, which will be a starting point for this article.
Theorem 3 (Theorem 1.1, [MR3885140]).
@@
The theorem above is sharp in the sense that for each prime $p\in\{2,3,5,7,11,13,17,37\}$ and each integer $n\in\mathcal{I}_{p}$, there exists a non-CM elliptic curve over $\mathbb{Q}$ admitting a rational isogeny of degree $p$ with index $n$. Such examples can be found using the search functionality of the LMFDB and are recorded in Appendix A. These values of $p$ are precisely the possible degrees of a prime degree rational isogeny for a non-CM elliptic curve over $\mathbb{Q}$ by Mazur’s theorem [MR482230]. The final sentence of the theorem follows as a consequence of the results for individual prime degrees. Indeed, if $E/\mathbb{Q}$ admits a rational cyclic isogeny of degree $N\geq 2$, then $E$ admits a rational isogeny of degree $p$ for each prime $p$ dividing $N$, so its adelic index belongs to $\mathcal{I}_{p}\subseteq\mathcal{I}$ for each such $p$.
@@
The proof of Theorem 4 extends throughout most of this paper. Section 2 reviews background on Galois representations, isogenies, twists, and modular curves. Section 3 discusses methods for computing rational points on modular curves. In Section 4, we determine all possible $\ell$-adic images containing I-I for non-CM elliptic curves admitting a rational isogeny of degree $p$ (and no rational isogeny of a larger prime degree) by studying rational points on modular curves and using Mazur and Kenku’s results on isogenies [MR482230, MR675184] along with Lemos’s theorem (Theorem 3). In Section 5, we use a result of Jones (Theorem 24) to give a product formula for the adelic index involving $\ell$-adic indices for certain small primes and an entanglement factor.
@@
We complete the proof of Theorem 4 in Sections 6, 7, and 8. In Section 6, we consider the primes $p\in\{11,17,37\}$, for which the proof is straightforward. In general, the adelic index of an elliptic curve over $\mathbb{Q}$ depends only on its $j$-invariant (Lemma 9), and in these cases, the modular curve $X_{0}(p)$ has only finitely many rational points. Thus, it suffices to compute the adelic index for one representative elliptic curve for each non-CM rational point on $X_{0}(p)$. In Section 7, we treat the primes $p\in\{5,7,13\}$ using a lattice-based approach that builds on ideas from [MR3500996, MR4468989, ZywinaOpen]. In principle, the same approach could be applied for $p\in\{2,3\}$, but it would be computationally impractical due to the number of groups involved. Instead, in Section 8 we take a product-based approach to handle the primes $p\in\{2,3\}$.
@@
This article makes extensive use of computations carried out in Magma, Version V2.28-21 [MR1484478]. The GitHub repository containing code for this article is:
@@
We make extensive use of data from the LMFDB [lmfdb]. Throughout the article, we use the LMFDB labeling conventions for elliptic curves and modular curves. For elliptic curves, the label N.aX specifies the conductor, followed by the isogeny class index and the isomorphism class index within the isogeny class. For modular curves, the label N.i.g.c.n specifies the level, index, genus, Gassmann class identifier, and an integer that distinguishes nonconjugate subgroups when the first four identifiers coincide. The modular curves section of the LMFDB is currently in beta. All LMFDB labels are hyperlinked to the object’s homepage.
@@ 1.1. Acknowledgments
We would like to thank Eran Assaf and Jeremy Rouse for their helpful suggestions related to computations of rational points on modular curves. The first-named author was partially supported by US-Israel BSF Grant 2022393.
@@ 2.1. Galois Representations
The $n$-adic Tate module of $E$ is the free $\mathbb{Z}_{n}$-module of rank $2$ defined by
@@
The natural action of $\operatorname{Gal}(\overline{F}/F)$ on $T_{n}(E)$ gives rise to the $n$-adic Galois representation
@@
We continue to let $E$ denote an elliptic curve over a number field $F$. Write $\chi\colon\operatorname{Gal}(\overline{F}/F)\to\widehat{\mathbb{Z}}^{\times}$ to denote the adelic cyclotomic character. It follows from the Weil pairing on $E$ that $\det\circ\rho_{E}=\chi$; see [MR2514094, §3.8] and [MR263823, IV-18 Lemma 2]. In particular, if $F=\mathbb{Q}$, then
@@
In other words, the composition $\det\circ\rho_{E}$ is surjective. Lastly, we note that in the case $F=\mathbb{Q}$, if $\ell\geq 5$ is a prime, then $\rho_{E,\ell}$ is surjective if and only if $\rho_{E,\ell^{\infty}}$ is surjective [MR263823, IV-23 Lemma 3].
@@ 2.2. Isogenies
In this subsection, we review some standard facts about isogenies and prove a lemma about the Galois images of elliptic curves related by a rational prime degree isogeny. First, let us recall the definition of an isogeny.
Definition 5.
Let $E$ and $E^{\prime}$ be elliptic curves over a number field $F$. An isogeny of elliptic curves
$\phi\colon E\longrightarrow E^{\prime}$
is a nonconstant morphism of curves that preserves base points. We call $\phi$ an $n$-isogeny if its kernel $C$ is a cyclic group of order $n$. We say that $\phi$ is $F$-rational if $C$ is stable under the action of $\operatorname{Gal}(\overline{F}/F)$, i.e., for all $\sigma\in\operatorname{Gal}(\overline{F}/F)$ and all $R\in C$, one has $\sigma(R)\in C$. Any cyclic subgroup $C\subseteq E(\overline{F})$ of order $n$ stabilized by $\operatorname{Gal}(\overline{F}/F)$ induces an $F$-rational $n$-isogeny to some elliptic curve $E^{\prime}/F$ whose kernel is $C$; see [MR2514094, Exercise 3.13]. If $F=\mathbb{Q}$, then instead of saying $\mathbb{Q}$-rational, we can simply say rational.
We now state a well-known result of Kenku, which extends Mazur’s [MR482230] classification of rational isogenies of prime degree to all degrees $n$ for which a rational $n$-isogeny exists.
Theorem 6 (Theorem 1, [MR675184]).
@@
In general, for an elliptic curve $E/F$ and a positive integer $n$, one has that $E$ admits an $F$-rational $n$-isogeny if and only if $G_{E}(n)$ is conjugate to a subgroup of the Borel subgroup
@@
Indeed, suppose that $E$ admits an $F$-rational $n$-isogeny $\theta\colon E\rightarrow E^{\prime}$. Then we can choose a point $P\in E[n]$ of order $n$ such that $\ker(\theta)=\langle P\rangle$. Since $\theta$ is $F$-rational, the subgroup $\langle P\rangle$ is also $F$-rational. Choose $Q$ in $E[n]$ such that $\{P,Q\}$ is a basis for $E[n]$. Because $\langle P\rangle$ is stable under the action of $\operatorname{Gal}(\overline{F}/F)$, we have for all $\sigma\in\operatorname{Gal}(\overline{F}/F)$,
$\sigma(P)=a_{\sigma}P,\hskip 14.22636pt\sigma(Q)=b_{\sigma}P+d_{\sigma}Q$
for some $a_{\sigma},b_{\sigma},d_{\sigma}\in\mathbb{Z}/n\mathbb{Z}$. Thus,
${\rho}_{E,n}(\sigma)=\begin{bmatrix}a_{\sigma}&b_{\sigma}\\ 0&d_{\sigma}\end{bmatrix}$.
This shows that $G_{E}(n)\subseteq B_{0}(n)$. To see why the converse is true, suppose now that $G_{E}(n)$ is conjugate to a subgroup of $B_{0}(n)$. Then we can choose a basis $\{P,Q\}$ such that $G_{E}(n)\subseteq B_{0}(n)$. Then $\langle P\rangle$ is stable under the action of $\operatorname{Gal}(\overline{F}/F)$, and hence is $F$-rational. It therefore induces an $F$-rational isogeny $E\rightarrow E^{\prime}$ with kernel $\langle P\rangle$ by [MR2514094, Exercise 3.13].
@@
Let $p$ be a prime number and $E/\mathbb{Q}$ be an elliptic curve. Then $E$ admits a rational $p$-isogeny if and only if $G_{E}(p^{\infty})$ is conjugate to a subgroup of $B_{0}(p^{\infty})$.
With our understanding of isogenies, we now state and prove a lemma relating the mod $n$ Galois images for isogenous elliptic curves (cf. [MR4868206, Corollary 2.21]).
@@ 2.3. Twists
Lemma 9 (Corollary 2.3, [zywina2022possibleindicesgaloisimage]).
@@ 2.5. Modular Curves
If $G\subseteq G^{\prime}$, then there is a natural morphism $X_{G}\to X_{G^{\prime}}$. In particular, associated to each modular curve $X_{G}$ is a morphism, called the $j$-map, $\pi_{G}\colon X_{G}\to X_{\operatorname{GL}_{2}(\widehat{\mathbb{Z}})}\simeq\mathbb{P}^{1}_{\mathbb{Q}}$. Then $\pi_{G}(X_{G}(\mathbb{Q}))\subseteq\mathbb{Q}\cup\{\infty\}$. The points of $X_{G}$ that lie above $\infty$ are known as cusps. For noncuspidal points, we have the following characterization. This result is well known (see, for instance, [zywina2015possibleimagesmodell, Proposition 3.3]).
@@
When $G$ is a Borel subgroup $B_{0}(N)$ of $\operatorname{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$, we denote $X_{G}$ by $X_{0}(N)$. By Proposition 14 and our discussion in Section 2.2, the rational noncuspidal points of $X_{0}(N)$ correspond to elliptic curves defined over $\mathbb{Q}$ (up to $\overline{\mathbb{Q}}$ isomorphism) that admit a rational $N$-isogeny. We write $X_{\mathrm{sp}}(N)$ and $X_{\mathrm{ns}}(N)$ to denote the modular curves associated with the split and nonsplit Cartan subgroups, respectively, and write $X_{\mathrm{sp}}^{+}(N)$ and $X_{\mathrm{ns}}^{+}(N)$ for the modular curves associated with their normalizers. Finally, we write $X_{S_{4}}(N)$ for the modular curve associated with the inverse image in $\operatorname{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ of the unique (up to conjugacy) subgroup of $\operatorname{PGL}_{2}(\mathbb{Z}/N\mathbb{Z})$ that is isomorphic to $S_{4}$. We will often also denote these modular curves by their LMFDB labels, as discussed in the introduction. With some abuse of notation, we will refer to $G$ and $X_{G}$ interchangeably when it is clear from the context.

### v2

@@ 1. Introduction
Theorem 1 (Serre’s Open Image Theorem, [40]).
@@
Serre’s Open Image Theorem is equivalent to the existence of a positive integer $n$ such that $G_{E}=\pi_{n}^{-1}(G_{E}(n))$; the least such $n$ is called the adelic level of $E$. Since Serre proved Theorem 1, there has been substantial progress toward understanding Galois images of non-CM elliptic curves over $\mathbb{Q}$. Notably, Zywina [44] and Sutherland [42] conjecturally classified all possible mod $\ell$ Galois images and gave efficient algorithms for computing them. Rouse–Zureick-Brown [37] and Rouse–Sutherland–Zureick-Brown [36] did similarly for $\ell$-adic images. In 2022, Zywina [46] gave a breakthrough algorithm for computing adelic images of elliptic curves.
An important invariant related to the adelic Galois image of a non-CM elliptic curve $E$ is the adelic index, the quantity $\operatorname{GL}_{2}(\widehat{\mathbb{Z}}):G_{E}]$. In general, the adelic index cannot be determined from the $\ell$-adic indices alone because of possible entanglements between the $\ell$-adic images. For example, consider the elliptic curve 37.a1 in the LMFDB. For all primes $\ell$, its $\ell$-adic image is $\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$, yet its adelic index is 2. In fact, the adelic index of any non-CM elliptic curve over $\mathbb{Q}$ must be at least $2$, as noted in [40, Proposition 22]. Moreover, Jones [22] proved that, when ordered by naive height, almost all elliptic curves over $\mathbb{Q}$ have adelic index $2$. Adelic indices greater than 2 occur for infinitely many $\overline{\mathbb{Q}}$-isomorphism classes [22, Theorem 4], but these are sparse asymptotically.
Zywina conjectured that there are only finitely many possible adelic indices for non-CM elliptic curves over $\mathbb{Q}$, and he even proposed a specific finite set $\mathcal{I}$ of all adelic indices [45, 46].
Conjecture 2 (Conjecture 1.5, [46]).
@@
This conjecture has been verified for all non-CM elliptic curves over $\mathbb{Q}$ currently appearing in the $L$-Functions and Modular Forms Database (LMFDB) [30], consisting of more than three million curves. The conjecture is closely connected to Serre’s uniformity question, which asks whether there exists a prime $p$ such that for any prime $\ell>p$, the mod $\ell$ Galois representation of any non-CM elliptic curve $E/\mathbb{Q}$ is surjective. It is now widely conjectured that Serre’s uniformity question has an affirmative answer with $p=37$.
Over the past several decades, significant progress toward Serre’s uniformity question has come from the study of rational points on modular curves along Dickson’s classification: if the mod $\ell$ representation is not surjective, its image is contained in a Borel subgroup, the normalizer of a split or nonsplit Cartan subgroup, or an exceptional subgroup. Serre [40] proved that the exceptional case cannot occur for $\ell>13$. Mazur [34] proved that if $\ell>37$, then $G_{E}(\ell)$ is not contained in a Borel subgroup. Bilu and Parent [7] proved that $G_{E}(\ell)$ cannot be contained in the normalizer of a split Cartan subgroup for sufficiently large $\ell$. Bilu, Parent, and Rebolledo [6] later showed that the same result holds for every prime $\ell\geq 11$, except possibly $\ell=13$. The latter case was subsequently resolved via quadratic Chabauty by Balakrishnan, Dogra, Müller, Tuitman, and Vonk [1]. As for the case of nonsplit Cartan subgroups, recent work of Furio and Lombardo [19] shows that if $\ell>37$ is nonsurjective, then $G_{E}(\ell)$ is precisely equal to the normalizer of a nonsplit Cartan subgroup. Showing that this does not happen for sufficiently large $\ell$ is the remaining problem for the uniformity question, but eliminating this case appears daunting.
Nevertheless, Serre’s uniformity question is known to have a positive answer for certain families of non-CM elliptic curves over $\mathbb{Q}$ due to work of Lemos. He proved that uniformity holds with $p=37$ for the family of non-CM elliptic curves $E/\mathbb{Q}$ for which there exists a prime $q$ such that $G_{E}(q)$ is contained in the normalizer of a split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ [29]. Further, he proved that uniformity holds with $p=37$ for the family of all non-CM elliptic curves admitting a nontrivial rational cyclic isogeny, which will be a starting point for this article.
Theorem 3 (Theorem 1.1, [28]).
@@
The theorem above is sharp in the sense that for each prime $p\in\{2,3,5,7,11,13,17,37\}$ and each integer $n\in\mathcal{I}_{p}$, there exists a non-CM elliptic curve over $\mathbb{Q}$ admitting a rational isogeny of degree $p$ with index $n$. Such examples can be found using the search functionality of the LMFDB and are recorded in Appendix A. These values of $p$ are precisely the possible degrees of a prime degree rational isogeny for a non-CM elliptic curve over $\mathbb{Q}$ by Mazur’s theorem [34]. The final sentence of the theorem follows as a consequence of the results for individual prime degrees. Indeed, if $E/\mathbb{Q}$ admits a rational cyclic isogeny of degree $N\geq 2$, then $E$ admits a rational isogeny of degree $p$ for each prime $p$ dividing $N$, so its adelic index belongs to $\mathcal{I}_{p}\subseteq\mathcal{I}$ for each such $p$.
@@
The proof of Theorem 4 extends throughout most of this paper. Section 2 reviews background on Galois representations, isogenies, twists, and modular curves. Section 3 discusses methods for computing rational points on modular curves. In Section 4, we determine all possible $\ell$-adic images containing I-I for non-CM elliptic curves admitting a rational isogeny of degree $p$ (and no rational isogeny of a larger prime degree) by studying rational points on modular curves and using Mazur and Kenku’s results on isogenies [34, 25] along with Lemos’s theorem (Theorem 3). In Section 5, we use a result of Jones (Theorem 23) to give a product formula for the adelic index involving $\ell$-adic indices for certain small primes and an entanglement factor.
@@
We complete the proof of Theorem 4 in Sections 6, 7, and 8. In Section 6, we consider the primes $p\in\{11,17,37\}$, for which the proof is straightforward. In general, the adelic index of an elliptic curve over $\mathbb{Q}$ depends only on its $j$-invariant (Lemma 8), and in these cases, the modular curve $X_{0}(p)$ has only finitely many rational points. Thus, it suffices to compute the adelic index for one representative elliptic curve for each non-CM rational point on $X_{0}(p)$. In Section 7, we treat the primes $p\in\{5,7,13\}$ using a lattice-based approach that builds on ideas from [37, 36, 46]. In principle, the same approach could be applied for $p\in\{2,3\}$, but it would be computationally impractical due to the number of groups involved. Instead, in Section 8 we take a product-based approach to handle the primes $p\in\{2,3\}$.
@@
This article makes extensive use of computations carried out in Magma, Version V2.28-21 [9]. The GitHub repository containing code for this article is:
@@
We make extensive use of data from the LMFDB [30]. Throughout the article, we use the LMFDB labeling conventions for elliptic curves and modular curves. For elliptic curves, the label N.aX specifies the conductor, followed by the isogeny class index and the isomorphism class index within the isogeny class. For modular curves, the label N.i.g.c.n specifies the level, index, genus, Gassmann class identifier, and an integer that distinguishes nonconjugate subgroups when the first four identifiers coincide. The modular curves section of the LMFDB is currently in beta. All LMFDB labels are hyperlinked to the object’s homepage.
@@ 1.1. Acknowledgments
We would like to thank Eran Assaf and Jeremy Rouse for their helpful suggestions related to computations of rational points on modular curves. We also thank Andrew Sutherland, Jeremy Rouse, and Jennifer Balakrishnan for their feedback on an earlier version of this article. The first-named author was partially supported by US-Israel BSF Grant 2022393.
@@ 2.1. Galois Representations
Each of the Galois images $G_{E},G_{E}(n^{\infty}),$ and $G_{E}(n)$ depends on the chosen basis and thus is well defined only up to conjugation. When defining the Galois representations, we adopt the convention that torsion points and Tate modules are represented as column vectors and that all matrix groups act by left multiplication. This agrees with the convention currently used in the LMFDB but differs from the convention used by Zywina in [46, 47]. Consequently, when using Zywina’s code to compute a Galois image or a modular curve, we must take a transpose.
@@
We continue to let $E$ denote an elliptic curve over a number field $F$. Write $\chi\colon\operatorname{Gal}(\overline{F}/F)\to\widehat{\mathbb{Z}}^{\times}$ to denote the adelic cyclotomic character. It follows from the Weil pairing on $E$ that $\det\circ\rho_{E}=\chi$; see [41, §3.8] and [39, IV-18 Lemma 2]. In particular, if $F=\mathbb{Q}$, then
@@
In other words, the composition $\det\circ\rho_{E}$ is surjective. Lastly, we note that in the case $F=\mathbb{Q}$, if $\ell\geq 5$ is a prime, then $\rho_{E,\ell}$ is surjective if and only if $\rho_{E,\ell^{\infty}}$ is surjective [39, IV-23 Lemma 3].
@@ 2.2. Isogenies
In this subsection, we review some standard facts about isogenies and prove a lemma about the Galois images of elliptic curves related by a rational prime degree isogeny.
First, let $\phi:E\rightarrow E^{\prime}$ be an isogeny of two elliptic curves over a number field $F$. We call $\phi$ an $n$-isogeny if its kernel $C$ is a cyclic group of order $n$. We say that $\phi$ is $F$-rational if $C$ is stable under the action of $\operatorname{Gal}(\overline{F}/F)$, i.e., for all $\sigma\in\operatorname{Gal}(\overline{F}/F)$ and all $R\in C$, one has $\sigma(R)\in C$. Any cyclic subgroup $C\subseteq E(\overline{F})$ of order $n$ stabilized by $\operatorname{Gal}(\overline{F}/F)$ induces an $F$-rational $n$-isogeny to some elliptic curve $E^{\prime}/F$ whose kernel is $C$; see [41, Exercise 3.13]. If $F=\mathbb{Q}$, then instead of saying $\mathbb{Q}$-rational, we simply say rational.
A well-known result of Kenku extends Mazur’s [34] classification of rational isogenies of prime degree to all degrees $n$ for which a rational $n$-isogeny exists.
Theorem 5 (Theorem 1, [25]).
@@
In general, for an elliptic curve $E/F$ and a positive integer $n$, one has that $E$ admits an $F$-rational $n$-isogeny if and only if $G_{E}(n)$ is conjugate to a subgroup of the Borel subgroup:
@@
Let $p$ be a prime number and $E/F$ be an elliptic curve. Then $E$ admits a rational $p$-isogeny if and only if $G_{E}(p^{\infty})$ is conjugate to a subgroup of $B_{0}(p^{\infty})$.
We now state and prove a lemma relating the mod $n$ Galois images for isogenous elliptic curves (cf. [13, Corollary 2.21]).
@@ 2.3. Twists
Lemma 8 (Corollary 2.3, [45]).
@@ 2.5. Modular Curves
We recall the following characterization of noncuspidal points. This result is well known (see [44, Proposition 3.3], for instance).
@@
When $G$ is a Borel subgroup $B_{0}(N)$ of $\operatorname{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$, we denote $X_{G}$ by $X_{0}(N)$. By Proposition 13 and our discussion in Section 2.2, the rational noncuspidal points of $X_{0}(N)$ correspond to elliptic curves defined over $\mathbb{Q}$ (up to $\overline{\mathbb{Q}}$ isomorphism) that admit a rational $N$-isogeny. We write $X_{\mathrm{sp}}(N)$ and $X_{\mathrm{ns}}(N)$ to denote the modular curves associated with the split and nonsplit Cartan subgroups, respectively, and write $X_{\mathrm{sp}}^{+}(N)$ and $X_{\mathrm{ns}}^{+}(N)$ for the modular curves associated with their normalizers. Finally, for $N$ prime, we write $X_{S_{4}}(N)$ for the modular curve associated with the inverse image in $\operatorname{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ of the unique (up to conjugacy) subgroup of $\operatorname{PGL}_{2}(\mathbb{Z}/N\mathbb{Z})$ that is isomorphic to $S_{4}$. We will often also denote these modular curves by their LMFDB labels, as discussed in the introduction. With some abuse of notation, we will refer to $G$ and $X_{G}$ interchangeably when it is clear from the context.